ABCD is a square and P,Q are the midpoints of BC,CD respectively.If AP=a and AQ=b,


1
Expert's answer
2021-02-02T03:39:39-0500


a = b


let x be the side of the square


"\vartriangle ADQ:"

AD = x; DQ = "\dfrac{x}{2}"; AQ = b = a


by the Pythagorean theorem:


"x^2 + \dfrac{x^2}{4} = a^2 = b^2"

"4x^2 + x^2 = 4a^2 = 4b^2"

"5x^2 = 4a^2 => x^2 = \dfrac{4a^2}{5} = \dfrac{4b^2}{5}" - area of a square ABCD